OpenStudy (anonymous):

PLEASE HELP GIVING MEDAL! Put in point-intercept form 5x+10y=40

TheSmartOne (thesmartone):

Do you mean slope intercept form.

OpenStudy (anonymous):

Yes! Sorry.

TheSmartOne (thesmartone):

Slope intercept form is y=mx+b

TheSmartOne (thesmartone):

so take the 5x to the other side

TheSmartOne (thesmartone):

It will become 10y=40-5x

TheSmartOne (thesmartone):

divide by 10 on both sides

TheSmartOne (thesmartone):

y=4-.5x

TheSmartOne (thesmartone):

Rearrange terms to make it in the format of y=mx+b

TheSmartOne (thesmartone):

y=-.5x+4

TheSmartOne (thesmartone):

Do you get what I did? Do you have any questions?

OpenStudy (anonymous):

Yes, I understand. Thankyou so much! I appreciate it. <3

TheSmartOne (thesmartone):

Your Welcome

OpenStudy (anonymous):

Could you help me with one more problem, not the same thing. It has to do with order of operations. I keep getting the wrong answer. :/

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Sure..

OpenStudy (anonymous):

Sorry, I don't want to seem like a bother. But, 47[3*8÷(3+1)]

OpenStudy (anonymous):

Ah, let me draw it out.

TheSmartOne (thesmartone):

Next time even if someone doesn't resond just post the question because someone else might click on the question and help you..

TheSmartOne (thesmartone):

If you see those question marks refresh the page and you can see it clearly...

OpenStudy (anonymous):

Oh okay, didn't know. Thanks! But, i've been trying to get the answer but neither are the correct answer.

TheSmartOne (thesmartone):

The Order of Operations PEMDAS P- Paranthesis E-Exponents M--Multiplication D-Division A-Addition S-Subtraction

OpenStudy (anonymous):

Yes, I know that. :/

TheSmartOne (thesmartone):

So first we see a big paranthesis and inside of it a smaller one. So lets take care of the smaller one.

TheSmartOne (thesmartone):

47[3*8÷(3+1)] 47[3*8÷(4)]

TheSmartOne (thesmartone):

Now we have to take care of the mutiplication inside of the paranthesis then the division inside the paranthesis.

OpenStudy (anonymous):

Yees.

TheSmartOne (thesmartone):

47[3*8÷(4)] 47[24÷4]

TheSmartOne (thesmartone):

47[24÷4] 47[6]

OpenStudy (anonymous):

So don't you multiply now?

TheSmartOne (thesmartone):

Now we have to do the multiplication 47[6]= 47(6)=282

TheSmartOne (thesmartone):

Did you get 282 last time...

TheSmartOne (thesmartone):

Next time you should show the steps you took so we can see where you went wrong and help you from there.

OpenStudy (anonymous):

Yes, and that's not one of the answers. There's 53, 56, 134, and 26. No 282, or even close to 282. :/

OpenStudy (anonymous):

So i'm really confused on why it wouldn't be on there?

OpenStudy (anonymous):

I did everything you did. But, I thought I got it wrong since the answer isn't on my homework. Would I just pick 134 since it's the highest..?

TheSmartOne (thesmartone):

Is it an online homework...

OpenStudy (anonymous):

Yeeep.

TheSmartOne (thesmartone):

Well maybe you can choose 134 but email your teacher about it...

OpenStudy (anonymous):

Thankyou for the help, I do appreciate you taking the time telling me how to get the answer. I sure will, thanks!

OpenStudy (anonymous):

Sorry for being a bother btw haha.

TheSmartOne (thesmartone):

No Problem..

TheSmartOne (thesmartone):

You were not a bother..